OpenStudy (anonymous):

Let sin(x)=3/5 and x is an acute angle. Find the exact value for sin (2x)

Parth (parthkohli):

Tons of ways to do this.\[\sin(2x) = 2\sin(x) \cos (x)\]Find the value of \(\cos(x) \) by using the identity \(\sin^2 x + \cos^2 x = 1\).

OpenStudy (anonymous):

Would cos(x) in this situation be 4/5 b/c of the Pythagorean theorem?

OpenStudy (anonymous):

Ahh I see now. 2(3/5)(4/5) sin2x=24/25. Thank you very much